# The contact process on dynamical random trees with degree dependence

### v1

@@ 7.3 Pushing the infection along a path
Given a single path of length $r$ on the vertices $\rho=x_{0},x_{1}\dots,x_{r}$ such that $\{x_{i-1},x_{i}\}\in\mathcal{E}$ for $1\leq i\leq r$, the goal in this subsection is to find a lower bound on the probability $\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}(x_{r}\in\mathbf{C}_{4r})$ where we let $IP^{\{\rho\}}_{\mathcal{T},N}(\cdot)=\mathbb{P}_{\mathcal{T},N}(\ \cdot\ |\ \mathbf{C}_{0}=\{\rho\})$. We begin by finding the probability to pass the infection along a single edge. In order to do so, we first make the following simple observation.
@@
For the second claim, we recall that the Laplace transform of the sum of two independent exponential random variables with parameter $a$ and $b$ has the form
$\frac{a}{a+\theta}\frac{b}{b+\theta}.$
By comparing factors, we see that $a$ and $b$ have to satisfy
$a+b=\alpha+\beta\quad\mbox{and}\quad ab=q\alpha\beta,$
which is true by taking $a$ and $b$ as in the statement of the lemma. ∎
@@
$a$ $:=\frac{\lambda+v(D_{x},D_{y})}{2}+\sqrt{\frac{1}{4}(\lambda+v(D_{x},D_{y}))^{2}-\lambda v(D_{x},D_{y})p(D_{x},D_{y})},$ (7.9)
$b:=\frac{\lambda+v(D_{x},D_{y})}{2}-\sqrt{\frac{1}{4}(\lambda+v(D_{x},D_{y}))^{2}-\lambda v(D_{x},D_{y})p(D_{x},D_{y})}.$
@@
For simplicity we drop the argument $(D_{x},D_{y})$ and write $v:=v(D_{x},D_{y})$ and $p=p(D_{x},D_{y})$. Suppose that initially the edge $\{x,y\}$ is closed and we denote by $\mathbb{P}(\,\cdot\,)$ referring to ($\mathbb{P}(\,\cdot\mid\{x,y\}\notin\mathbf{B}_{0})$. Set $S_{0}^{\rm cl/inf}=0$ and for $i\geq 1$, iteratively define
@@
Finally we will show a lower bound for the probability of pushing the infection along a path of bounded degree to a stable star. The proof uses a similar line of arguments as in [4, Lemma 5.1], however the presence of the dynamical graph structure leads to some changes. In order to do so, we first introduce some notation.
@@
For any vertex $x\in\mathcal{V}_{r}$, we define the subgraph $\mathcal{T}^{\rho-x}$ where the vertex set is given through $\mathcal{V}^{\rho-x}:=\{\rho,x_{1},x_{2},\dots,x_{r-1},x\}$ and we consider all edges $\{x,y\}\in\mathcal{E}$ such that $x,y\in\mathcal{V}^{\rho-x}$. Now we introduce the process $(\mathbf{C}_{t}^{\rho-x})_{t\geq 0}$, which is defined via the same graphical representation as $(\mathbf{C}_{t})_{t\geq 0}$ but the infection is restricted to $\mathcal{T}^{\rho-x}$, i.e.
@@
This means that only vertices contained in $\mathcal{V}^{\rho-x}$ participate in infection events, and therefore, as long as $\mathbf{C}_{0}\subset\mathcal{V}^{\rho-x}$ it follows that $\mathbf{C}^{\rho-x}_{t}\subset\mathcal{V}^{\rho-x}$ for all $t\geq 0$.
Now we define $\mathbf{P}_{r}$ as the set of all vertices in generation $r$ with degree $N+1$ which are connected to $\rho$ via a path consiting of vertices with degree bounded by $L$, i.e.
@@
We emphasize here that the notation $\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}(\cdot)$ in what follows corresponds to the conditional probability on the tree and on the event that the root $\rho$ is infected and the background process starts in any distribution.
@@
Let $r\in\mathbb{N}$. Suppose the background process is started in an arbitrary distribution. Let $x\in\mathcal{V}_{r}$ and $\rho=x_{0},x_{1},\dots,x_{r}=x$ is the unique path from $\rho$ to $x$ then there exists a constant $\gamma>0$ independent of $x$ such that
@@
Moreover, let $N,L\in\mathbb{N}$ with $N\geq L$ and suppose that $v(\cdot,\cdot)$ and $p(\cdot,\cdot)$ satisfy (2.3). Then for $x\in\mathbf{P}_{r}$ it follows that
@@
Furthermore, define $\tau:=\sum_{i=1}^{r}t_{i}$ with $t_{i}=s_{i}-s_{i-1}$ for $i\in\{1,\dots,r\}$. Also denote by
@@
$\mathbb{P}^{\{\rho\}}_{\mathcal{T}}\Big{(}x_{r}\in\bigcup_{s\leq 4r}\mathbf{C}^{\rho-x}_{s}\Big{)}\geq\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(B\cap\{\tau\leq 4r\})=\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\tau\leq 4r\mid B)\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(B).$ (7.15)
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(B_{i}\ |\ \{x_{i-1},x_{i}\}\not\in\textbf{B}_{s_{i-1}})=\frac{\lambda v(D_{x_{i-1}},D_{x_{i}})p(D_{x_{i-1}},D_{x_{i}})}{\lambda+v(D_{x_{i-1}},D_{x_{i}})+\lambda v(D_{x_{i-1}},D_{x_{i}})p(D_{x_{i-1}},D_{x_{i}})+1}.$
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(B_{i})\geq\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(B_{i}\ |\ \{x_{i-1},x_{i}\}\not\in\textbf{B}_{s_{i-1}}),$
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(B)$ $\geq\prod_{i=1}^{r}\mathbb{P}^{\{\rho\}}(B_{i}\ |\ \{x_{i-1},x_{i}\}\not\in\textbf{B}_{s_{i-1}})$
@@
On the other hand, appealing to Markov’s inequality and the definition of $B$ and $\tau$, we have, for any $\theta>0$
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\tau\geq 4r\ |\ B)$ $=\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(e^{\theta\tau}\geq e^{4\theta r}\ |\ B)\leq e^{-4\theta r}\mathbb{E}_{\mathcal{T}}^{\{\rho\}}\left[e^{\theta\tau}\ \Big{|}\Big{.}\ B\right]$
$\leq e^{-4\theta r}\mathbb{E}_{\mathcal{T}}^{\{\rho\}}\bigg{[}\prod_{i=1}^{r}e^{\theta t_{i}}\ \Big{|}\Big{.}\ \bigcap_{i=1}^{r}B_{i}\cap\{\{x_{i-1},x_{i}\}\not\in\textbf{B}_{s_{i-1}}\}\bigg{]}$
$\leq e^{-4\theta r}\prod_{i=1}^{r}\mathbb{E}_{\mathcal{T}}^{\{\rho\}}\Big{[}e^{\theta t_{i}}\ \Big{|}\Big{.}\ B_{i}\cap\{\{x_{i-1},x_{i}\}\not\in\textbf{B}_{s_{i-1}}\}\Big{]},$
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(t_{i}>s\ |\ B_{i}\cap\{\{x_{i-1},x_{i}\}\not\in\textbf{B}_{s_{i-1}}\})=\frac{b_{i}+1}{b_{i}-a_{i}}e^{-(a_{i}+1)s}+\frac{a_{i}+1}{a_{i}-b_{i}}e^{-(b_{i}+1)s},$
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}\left(\tau\geq 4r\mid B\right)$ $\leq e^{-4\theta r+r\log\phi(\theta)},$
where $\phi(\theta):=\mathbb{E}_{\mathcal{T}}[e^{\theta r_{i}}]=(1-\theta)^{-2}$. Now, note that
$\lim_{\theta\to 0}\frac{\log(1-\theta)^{-2}}{\theta}=2.$
Therefore, by choosing $\theta>0$ small enough, we can deduce that there exists $\gamma>0$ such that
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\tau\geq 4r\ |\ B)\leq e^{-r\gamma}.$ (7.17)
Plugging (7.3) and (7.17) back into (7.15), we get the first statement. For the second claim, we assume that the update speed $v$ and $p$ satisfy (2.3), and that $x\in\mathbf{P}_{r}$. First, we deal with the case $\eta\geq 0$. Since $D_{x_{i}}\geq 1$ for all $0\leq i\leq r$ in this case we have $v(\cdot,\cdot)\geq\nu_{1}$, which yields
@@
$\frac{\lambda v(D_{x_{i-1}},D_{x_{i}})p(D_{x_{i-1}},D_{x_{i}})}{\lambda+v_{\eta}(D_{x_{i-1}},D_{x_{i}})+\lambda v(D_{x_{i-1}},D_{x_{i}})p(D_{x_{i-1}},D_{x_{i}})+1}\geq p(D_{x_{i-1}},D_{x_{i}})\frac{\lambda\nu_{1}}{\lambda+\lambda\nu_{1}+\nu_{1}+1}.$
@@ 7.4 Transmitting the infection from one star to another
This means that only vertices contained in $\mathcal{V}^{*-x}$ participate in infection events, and therefore, as long as $\mathbf{C}_{0}\subset\mathcal{V}^{*-x}$ it follows that $\mathbf{C}^{*-x}_{t}\subset\mathcal{V}^{*-x}$ for all $t\geq 0$.
If we start with only the centre of a star infected, we estimate the probability to push the infection to the next star. We use Lemmas 7.3 and 7.7 to find a lower bound for the probability that the star $x$, which is at distance $r$ from the root and has the same degree as the root, is infected before the infection dies out around the root.
Before we proceed, we recall some notation. We defined the set of good infected neighbours by $\mathcal{G}_{k}^{\prime}(z)$ in (7.6), where a given $z\in\mathcal{N}_{\rho}$ is excluded. Similar we defined the the $\sigma$-algebra $\mathcal{F}_{s,t}(z)$ to be generated by the graphical construction up to time $s$, and the processes $\Delta^{\text{rec}}_{y}$ and $\Delta^{\text{up}}_{\{\rho,y\}}$ for all $y\in\mathcal{N}_{\rho}\backslash\{z\}$ up to time $t$. Note that for the sake of readability we omit $z$, whenever it is clear from the context. Lastly at the beginning of Subsection 7.2 we defined $T=T_{N}=(1+v_{2}N^{\eta})^{-1}$.
@@
We set initially $X_{0}=0$. Now at any time point in $\Delta^{M,\inf}$ the process $X$ jumps to state $1$ and at any time point in $\Delta^{\text{rec}}_{\rho}$ it jumps into state 00, i.e. we get a Markov process with transitions
[…]
@@
(see e.g. [16, Example 2.6]). Using the inequality $1-e^{-x}\geq\frac{x}{x+1}$ which holds for all $x>-1$, we see that
@@
The next lemma provides an upper bound for the probability that the number of infected good neighbours during the time interval $[0,(\bar{k}+2)T]$ is at most $\delta\lambda TNp(N,L)$ given that the root is initially infected and has degree $N$, where $\bar{k}=\lfloor e^{\delta\lambda^{2}T^{2}Np(N,L)/4}\rfloor$ as defined in Lemma 7.3 for $\delta=c_{L}/8$ with $c_{L}$ chosen as in Lemma 7.1. Finally, recall that $\mathcal{S}(z)=\mathcal{S}_{\rho,N}(z)$ with $z\in\mathcal{N}$ is the event that $\rho$ is a stable star, where $z$ is excluded, defined in (7.5).
@@
$\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\Big{(}\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{k}^{\prime}(z)|\leq\delta\lambda TNp(N,L)\,\Big{|}\,\mathcal{F}_{0,S}(z)\Big{)}\leq R_{\lambda}(N),$
on the event $\mathcal{S}(z)$, where
@@
In the whole proof we exclude $z\in\mathcal{N}$ from the star. Thus, we suppress it in the notation and only write $\mathcal{S}$, $\mathcal{G}_{k}^{\prime}$ and $\mathcal{F}_{0,S}$. We begin by observing
$\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}\Big{(}$ $\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{k}^{\prime}|\geq\delta\lambda TNp(N,L)\ \Big{|}\ \mathcal{F}_{0,S}\Big{)}$
@@
holds on the event $\mathcal{S}$. In summary we get that
$\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}\Big{(}\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{k}^{\prime}|$ $\leq\delta\lambda TNp(N,L)\big{|}\ \mathcal{F}_{0,S}\Big{)}$
$\leq 1-(1-Ce^{-\delta\lambda^{2}T^{2}Np(N,L)})e^{-2T}(1-e^{-\delta\lambda TNp(N,L)})$
holds on the event $\mathcal{S}$, which proves the claim. ∎
@@
$\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}\left(x\not\in\mathbf{C}^{*-x}_{s}\ \text{for all}\ s\in[0,t]\ |\ \mathcal{F}_{0,S}(x_{1})\right)\leq F_{\lambda}(N,r)+R_{\lambda}(N),$
for all $t\in[\tfrac{S}{2},S]$ on the event $\mathcal{S}_{\rho,N}(x_{1})$, where $S$ is defined as in (7.18), $R_{\lambda}(N)$ as in (7.19),
$F_{\lambda}(N,r):=\Bigg{(}1-b_{\lambda}(N)C_{p}(N)\bigg{(}\frac{\lambda\nu_{1}\kappa_{1}c_{p}}{\lambda+\lambda\nu_{1}+\nu_{1}+1}\bigg{)}^{r}\Bigg{)}^{\left\lfloor S(8r+4T)^{-1}\right\rfloor}$ (7.20)
@@
For the sake of readability we will again omit if possible the subscripts $\rho$ and the argument $x_{1}$ and write $\mathcal{S}$, $\mathcal{G}^{\prime}_{k}$ and $\mathcal{F}_{0,S}$ throughout. The function
$t\mapsto\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}\left(x\not\in\mathbf{C}^{*-x}_{s}\ \text{for all}\ s\in[0,t]\ |\ \mathcal{F}_{0,S}(x_{1})\right)$
is monotone decreasing, and therefore it suffices to derive the bound for $t=\frac{S}{2}$. Let $r\geq 1$ and set $m:=\lfloor S(8r+4T)^{-1}\rfloor$. Recall that we denote the unique path from $\rho$ to $x$ by $\rho=x_{0},x_{1}\dots,x_{r}=x$. We begin by noting that
$\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}\left(x\not\in\mathbf{C}_{s}^{*-x}\ \text{for all}\ s\in[0,\tfrac{S}{2}]\ |\ \mathcal{F}_{0,S}\right)$
$\hskip 2.84526pt\leq\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}\Big{(}\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{k}^{\prime}|\leq\delta\lambda TNp(N,L)\ |\ \mathcal{F}_{0,S}\Big{)}$ (7.22)
@@
Thanks to Lemma 7.9 we know that the first probability on the right-hand side is bounded from above by $R_{\lambda}(N)$ on the event $\mathcal{S}$.
[…]
@@
where we used the tower property, since $\mathcal{F}_{t_{m},t^{*}_{m}+2T,S}\subset\mathcal{F}_{t^{*}_{m}+2T,S}$, and that $B^{1}_{m}\in\mathcal{F}_{t^{*}_{m}+2T,S}$ by definition. Recall that before Lemma 7.7 we defined the infection process $\mathbf{C}^{\rho-x}$, which is restricted to the path from $\rho$ to $x$. By definition it follows that $\mathbf{C}^{\rho-x}_{t}\subset\mathbf{C}^{*-x}_{t}$ for all $t\geq 0$. Let us empasize at this point that by excluding $x_{1}$ from the set of good neighbours the path leading from $\rho$ to $x$ and the set $\mathcal{G}^{\prime}_{k}(x_{1})$ are defined on disjoint parts of the graphical representation for all $k$. This enables us to use the Markov property and monotonicity to obtain that
@@
Moreover, appealing to Lemmas 7.7 and 7.8 we get

### v2

@@ 7.3 Pushing the infection along a path
Given a single path of length $r$ on the vertices $\rho=x_{0},x_{1}\dots,x_{r}$ such that $\{x_{i-1},x_{i}\}\in\mathcal{E}$ for $1\leq i\leq r$, the goal in this subsection is to find a lower bound on the probability $\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}(x_{r}\in\mathbf{C}_{4r})$ where we let $\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}(\cdot)=\mathbb{P}_{\mathcal{T},N}(\ \cdot\ |\ \mathbf{C}_{0}=\{\rho\})$. We begin by finding the probability to pass the infection along a single edge. In order to do so, we first make the following simple observation.
@@
For the second claim, we recall that the Laplace transform of the sum of two independent exponential random variables with parameter $a$ and $b$ has the form $a(a+\theta)^{-1}b(b+\theta)^{-1}$. By comparing factors, we see that $a$ and $b$ have to satisfy $a+b=\alpha+\beta$ and $ab=q\alpha\beta,$ which is true by taking $a$ and $b$ as in the statement of the lemma. ∎
@@
$a$ $=\frac{\lambda+v(D_{x},D_{y})}{2}+\sqrt{\frac{1}{4}(\lambda+v(D_{x},D_{y}))^{2}-\lambda v(D_{x},D_{y})p(D_{x},D_{y})},$ (7.8)
$b=\frac{\lambda+v(D_{x},D_{y})}{2}-\sqrt{\frac{1}{4}(\lambda+v(D_{x},D_{y}))^{2}-\lambda v(D_{x},D_{y})p(D_{x},D_{y})}.$
@@
For simplicity we drop the argument $(D_{x},D_{y})$ and write $v:=v(D_{x},D_{y})$ and $p=p(D_{x},D_{y})$. Suppose that initially the edge $\{x,y\}$ is closed and we denote by $\mathbb{P}(\,\cdot\,)$ referring to $\mathbb{P}(\,\cdot\mid\{x,y\}\notin\mathbf{B}_{0})$. Set $S_{0}^{\rm cl/inf}=0$ and for $i\geq 1$, iteratively define
@@
Finally we will show a lower bound for the probability of pushing the infection along a path of bounded degree to a stable star. The proof uses a similar line of arguments as in [5, Lemma 5.1], however the presence of the dynamical graph structure leads to some changes. In order to do so, we first introduce some notation.
@@
For any vertex $x\in\mathcal{V}_{r}$, we define the subgraph $\mathcal{T}^{\rho-x}$ where the vertex set is given through $\mathcal{V}^{\rho-x}:=\{\rho,x_{1},\dots,x_{r-1},x\}$ and we consider all edges $\{x,y\}\in\mathcal{E}$ such that $x,y\in\mathcal{V}^{\rho-x}$. Now we introduce the process $(\mathbf{C}_{t}^{\rho-x})_{t\geq 0}$, which is defined via the same graphical representation as $(\mathbf{C}_{t})_{t\geq 0}$ but the infection is restricted to $\mathcal{T}^{\rho-x}$, i.e.
@@
This means that only vertices contained in $\mathcal{V}^{\rho-x}$ participate in infection events, and therefore $\mathbf{C}^{\rho-x}_{t}\subset\mathcal{V}^{\rho-x}$ for all $t\geq 0$.
Now we define $\mathbf{P}_{r}$ as the set of all vertices in generation $r$ with degree $N+1$ which are connected to $\rho$ via a path consisting of vertices with degree bounded by $L$, i.e.
@@
[…]
@@
Let $r\in\mathbb{N}$. Suppose the background process is started in an arbitrary distribution. Let $x\in\mathcal{V}_{r}$ and $\rho=x_{0},x_{1},\dots,x_{r}=x$ be the unique path from $\rho$ to $x$ then there exists a constant $\gamma>0$ independent of $x$ and $r$ such that
@@
Moreover, let $N,L\in\mathbb{N}$ with $N\geq L$ and suppose that $p(\cdot,\cdot)$ and $v(\cdot,\cdot)$ satisfy (2.3) and (2.4), respectively. Then for $x\in\mathbf{P}_{r}$ it follows that
@@
Furthermore, define $t_{i}=s_{i}-s_{i-1}$ for $i\in\{1,\dots,r\}$. Also denote by
@@
$\mathbb{P}^{\{\rho\}}_{\mathcal{T}}\Big(x_{r}\in\bigcup_{s\leq 4r}\mathbf{C}^{\rho-x}_{s}\Big)\geq\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(B\cap\{s_{r}\leq 4r\})=\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(s_{r}\leq 4r\mid B)\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(B).$ (7.14)
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(B_{i}\ |\ \{x_{i-1},x_{i}\}\not\in\mathbf{B}_{s_{i-1}})=\frac{\lambda v(D_{x_{i-1}},D_{x_{i}})p(D_{x_{i-1}},D_{x_{i}})}{\lambda+v(D_{x_{i-1}},D_{x_{i}})+\lambda v(D_{x_{i-1}},D_{x_{i}})p(D_{x_{i-1}},D_{x_{i}})+1}.$
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(B_{i})\geq\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(B_{i}\ |\ \{x_{i-1},x_{i}\}\not\in\mathbf{B}_{s_{i-1}}),$
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(B)$ $\geq\prod_{i=1}^{r}\mathbb{P}^{\{\rho\}}(B_{i}\ |\ \{x_{i-1},x_{i}\}\not\in\mathbf{B}_{s_{i-1}})$
@@
On the other hand, appealing to Markov’s inequality and the definition of $B$ and $s_{r}$, we have, for any $\theta>0$
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(s_{r}\geq 4r\ |\ B)$ $=\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(e^{\theta s_{r}}\geq e^{4\theta r}\ |\ B)\leq e^{-4\theta r}\mathbb{E}_{\mathcal{T}}^{\{\rho\}}\left[e^{\theta s_{r}}\ \Big|\Big.\ B\right]$
$\leq e^{-4\theta r}\mathbb{E}_{\mathcal{T}}^{\{\rho\}}\bigg[\prod_{i=1}^{r}e^{\theta t_{i}}\ \Big|\Big.\ \bigcap_{i=1}^{r}B_{i}\cap\{\{x_{i-1},x_{i}\}\not\in\mathbf{B}_{s_{i-1}}\}\bigg]$
$\leq e^{-4\theta r}\prod_{i=1}^{r}\mathbb{E}_{\mathcal{T}}^{\{\rho\}}\Big[e^{\theta t_{i}}\ \Big|\Big.\ B_{i}\cap\{\{x_{i-1},x_{i}\}\not\in\mathbf{B}_{s_{i-1}}\}\Big],$
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(t_{i}>s\ |\ B_{i}\cap\{\{x_{i-1},x_{i}\}\not\in\mathbf{B}_{s_{i-1}}\})=\frac{b_{i}+1}{b_{i}-a_{i}}e^{-(a_{i}+1)s}+\frac{a_{i}+1}{a_{i}-b_{i}}e^{-(b_{i}+1)s},$
@@
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}\left(s_{r}\geq 4r\mid B\right)$ $\leq e^{-4\theta r+r\log\phi(\theta)},$
where $\phi(\theta):=\mathbb{E}_{\mathcal{T}}[e^{\theta r_{i}}]=(1-\theta)^{-2}$. Now, note that $\lim_{\theta\to 0}\frac{\log((1-\theta)^{-2})}{\theta}=2.$ Therefore, by choosing $\theta>0$ small enough, we can deduce that there exists $\gamma>0$ (independent of $r$) such that
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(s_{r}\geq 4r\ |\ B)\leq e^{-r\gamma}.$ (7.16)
Plugging (7.3) and (7.16) back into (7.14), we get the first statement. For the second claim, we assume that $p(\cdot,\cdot)$ and $v(\cdot,\cdot)$ satisfy (2.3) and (2.4), respectively, and that $x\in\mathbf{P}_{r}$. First, we deal with the case $\eta\geq 0$. Since $D_{x_{i}}\geq 1$ for all $0\leq i\leq r$ in this case we have $v(\cdot,\cdot)\geq\nu_{1}$ and using the fact that the mapping $v\mapsto\lambda vp(\lambda+v+\lambda vp+1)^{-1}$ is increasing, we deduce
@@
$\frac{\lambda v(D_{x_{i-1}},D_{x_{i}})p(D_{x_{i-1}},D_{x_{i}})}{\lambda+v(D_{x_{i-1}},D_{x_{i}})+\lambda v(D_{x_{i-1}},D_{x_{i}})p(D_{x_{i-1}},D_{x_{i}})+1}\geq p(D_{x_{i-1}},D_{x_{i}})\frac{\lambda\nu_{1}}{\lambda+\lambda\nu_{1}+\nu_{1}+1}.$
@@ 7.4 Transmitting the infection from one star to another
This means that only vertices contained in $\mathcal{V}^{*-x}$ participate in infection events, and therefore $\mathbf{C}^{*-x}_{t}\subset\mathcal{V}^{*-x}$ for all $t\geq 0$.
If we start with only the centre of a star infected, we estimate the probability to push the infection to the next star. We use Lemmas 7.2 and 7.6 to find a lower bound for the probability that the star $x$, which is at distance $r$ from the root and has the same degree as the root, is infected before the infection dies out around the root.
Before we proceed, we recall some notation. We defined the set of good infected neighbours by $\mathcal{G}_{k}^{\prime}$ in (7.5), $\mathcal{F}_{s,t}$ the $\sigma$-algebra generated by the graphical construction up to time $s$, and the processes $\Delta^{\text{rec}}_{y}$ and $\Delta^{\text{up}}_{\{\rho,y\}}$ for all $y\in\mathcal{N}_{\rho}$ up to time $t$. Lastly at the beginning of Subsection 7.1 we defined $T=T_{N}=(1+v_{2}N^{\eta})^{-1}$.
@@
We set initially $X_{0}=0$. Now at any time point in $\Delta^{M,\inf}$ the process $X$ jumps to state $1$ and at any time point in $\Delta^{\text{rec}}_{\rho}$ it jumps into state 0, i.e. we get a Markov process with transitions
$0\to 1$  $\quad\text{ with rate }\lambda M$
$1\to 0$  $\quad\text{ with rate }1.$
@@
(see e.g. [16, Example 2.6]). Using the inequality $1-e^{-x}\geq\frac{x}{x+1}$ which holds for all x>1x>-1, we see that
@@
The next lemma provides an upper bound for the probability that, at some point in the time interval $0,(\bar{k}+2)T]$, the number of infected good neighbors is less than or equal to $\delta\lambda TNp(N,L)$ given that the root is initially infected and has degree $N$, where $\bar{k}=\lfloor e^{\delta\lambda^{2}T^{2}Np(N,L)/4}\rfloor$ as defined in Lemma 7.2 for $\delta=c_{L}/8$ with $c_{L}$ chosen as in Lemma 7.1.
@@
$\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\Big(\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{k}^{\prime}|\leq\delta\lambda TNp(N,L)\,\Big|\,\mathcal{F}_{0,S}\Big)\leq R_{\lambda}(N),$
on the event $\mathcal{S}$ (defined in (7.2)), where
@@
We begin by observing
[…]
@@
holds on the event $\mathcal{S}$. This completes the proof. ∎
Remark 7.9.
For technical reasons, we need to ensure that the reservoir around $\rho$ and the path towards the next star share only the centre $\rho$, and are otherwise defined on disjoint parts of the graph. This technicality poses no real issue, since it is not difficult to realise that all the results in Section 7.1, as well as Lemmas 7.7 and 7.8, still hold when a single neighbour $x_{1}\in\mathcal{N}_{\rho}$ is excluded. To avoid introducing additional notation, we continue to use the symbols $\mathcal{G}_{\rho,k}$, $\mathcal{G}^{\prime}_{\rho,k}$, $\mathcal{F}_{s,t}$, and $\mathcal{S}_{\rho,N}$ throughout the remainder of the paper, with the convention that these objects are now defined excluding the vertex $x_{1}\in\mathcal{N}_{\rho}$.
@@
$\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}\left(x\not\in\mathbf{C}^{*-x}_{s}\ \text{for all}\ s\in[0,\tfrac{S}{2}]\ |\ \mathcal{F}_{0,S}\right)\leq F_{\lambda}(N,r)+R_{\lambda}(N),$
on the event $\mathcal{S}_{\rho,N}$, where $S$ is defined as in (7.17), $R_{\lambda}(N)$ as in (7.18),
$F_{\lambda}(N,r)=\Bigg(1-b_{\lambda}(N)C_{p}(N)\bigg(\frac{\lambda\nu_{1}\kappa_{1}c_{p}}{\lambda+\lambda\nu_{1}+\nu_{1}+1}\bigg)^{r}\Bigg)^{\left\lfloor S(8r+4T)^{-1}\right\rfloor}$ (7.19)
@@
Let $r\geq 1$ and set $m:=\lfloor S(8r+4T)^{-1}\rfloor$. Recall that we denote the unique path from $\rho$ to $x$ by $\rho=x_{0},x_{1},\dots,x_{r}=x$. We begin by noting that
$\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}\left(x\not\in\mathbf{C}_{s}^{*-x}\ \text{for all}\ s\in[0,\tfrac{S}{2}]\ |\ \mathcal{F}_{0,S}\right)\leq\mathbb{P}^{\{\rho\}}_{\mathcal{T},N}\Big(\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{k}^{\prime}|\leq\delta\lambda TNp(N,L)\ |\ \mathcal{F}_{0,S}\Big)$ (7.21)
@@
Thanks to Lemma 7.8 and Remark 7.9 we know that the first probability on the right-hand side is bounded from above by $R_{\lambda}(N)$ on the event $\mathcal{S}$. Thus it is enough to derive the upper bound for the second probability. Define the sequence of times $\{t_{i},\ 1\leq i\leq m+1\}$ such that $t_{1}=0$ and $t_{i+1}-t_{i}=(4r+2T)$ for $i\in\{1,\dots,m\}$. Now set $D_{i}:=\big\{|\mathcal{G}_{\lfloor t_{i}/T\rfloor}^{\prime}(x_{1})|\geq\delta\lambda TNp(N,L)\}$ for $i\in\{2,\dots,m\}$, then we see that
@@
where we used the tower property, since $\mathcal{F}_{t_{m},t^{*}_{m}+2T,S}\subset\mathcal{F}_{t^{*}_{m}+2T,S}$, and that $B^{1}_{m}\in\mathcal{F}_{t^{*}_{m}+2T,S}$ by definition. Recall that before Lemma 7.6 we defined the infection process $\mathbf{C}^{\rho-x}$, which is restricted to the path from $\rho$ to $x$. By definition it follows that $\mathbf{C}^{\rho-x}_{t}\subset\mathbf{C}^{*-x}_{t}$ for all $t\geq 0$. Let us emphasize at this point that by excluding $x_{1}$ from the set of good neighbours the path leading from $\rho$ to $x$ and the set $\mathcal{G}^{\prime}_{k}$ are defined on disjoint parts of the graphical representation for all $k$ (see Remark 7.9). This enables us to use the Markov property and monotonicity to obtain that
@@
Moreover, appealing to Lemmas 7.6 and 7.7 we get